Use the given claim and test statistic to find the P - value.

Claim: The mean monthly salary of employees at one company is at most $32,500.
Test statistic: z = 0.52

0.6985

0.6030

0.3015

0.3970
Answer #1

Solution :

Given that,

This is the right tailed test ,

z = 0.52

P(z > 0.52) = 1 - P(z < 0.52) = 1 - 0.6985 = 0.3015

P-value = 0.3015